1st one is....
"A coin bank contains nickels,dimes and quarters totaling $5.45. If there are twice as many quarters as dimes, and 11 more nickels than quarters, how many of each coin are in the bank?"
2nd one:
"A vending machine contains 6 times as many quarters as dimes. If the total amount of money in the machine is $28.50, how many quarters are there?
and the last one:
"A solution of 63% is mixed with a solution of 43% to form 40 liters of a 53% solution. How much of the 63% solution must she use? (In liters)"
Thanks :)

Answer by mananth(16946)

1st one is....
"A coin bank contains nickels,dimes and quarters totaling $5.45. If there are twice as many quarters as dimes, and 11 more nickels than quarters, how many of each coin are in the bank?"
let dimes = x
quarters = 2x
nickels = 2x+11
Total money = 545 cents
10x+25*2x+5(2x+11)=545
10x+50x+10x+55=545
70x=545-55
70x=490
x=7 dimes
2x= 2*7= 14 quarters
nickels = 2x+11=2*7+11=25
..
2nd one:
"A vending machine contains 6 times as many quarters as dimes. If the total amount of money in the machine is $28.80, how many quarters are there?
dimes = x
quarters = 6x
money 2880 cents
..
10x+25*6x=2800
10x+150x=2880
160x=2880
x= 18 dimes
6x = 108 quarters
..

and the last one:
"A solution of 63% is mixed with a solution of 43% to form 40 liters of a 53% solution. How much of the 63% solution must she use? (In liters)"
63% =x liters
43%= 40-x liters
53% 40 liters
..
0.63x+0.43(40-x)= 0.53*40
0.63x+17.20-0.43x= 21.2
0.2x = 4
x=20 liters